OpenStudy (anonymous):

Number Theory. Deduce whether : 5^ (12) + 2^(10) is prime or not.

OpenStudy (anonymous):

2^10 =1024, easy to figure out 5^12 is going to end in a 0 or 5 hmm... if it ends in a 5 your last digit is going to be seven...

OpenStudy (anonymous):

doesn't help much as 19 is prime and 39 is not

OpenStudy (anonymous):

so your looking at xxx...xxx(24+25)

OpenStudy (anonymous):

14657×16657 (2 distinct prime factors) = 244141649 Anybody want to check my arithmetic?

OpenStudy (anonymous):

nope I trust wolfram...

OpenStudy (anonymous):

wolfam will factor in to the product of two primes, but i am fairly sure that is not the point of the problem

OpenStudy (anonymous):

In short, no, it's not a prime # if it has factors.

OpenStudy (anonymous):

yes... but how do you figure that out...

OpenStudy (anonymous):

Prime factorization is what I used

OpenStudy (anonymous):

an even power of 5 will always end in 625 (if power is greater than 2, obviously) if it's any help

OpenStudy (anonymous):

5*5*5*5*5*5*5*5*5*5*5*5*2*2*2*2*2*2*2*2*2*2 (hopefully I counted right)

OpenStudy (anonymous):

I don't know if this theory is correct by somebody told me a long time ago that even powers of base 2 never result in prime #'s

OpenStudy (turingtest):

*

OpenStudy (anonymous):

but somebody* ack typos

OpenStudy (anonymous):

Anybody willing to disprove that theory? :D

OpenStudy (anonymous):

any power of anything will never be prime, unless prime doesn't mean what I think it means?

OpenStudy (anonymous):

5*5*5*5*5*5*5*5*5*5*5*5+2*2*2*2*2*2*2*2*2*2

OpenStudy (anonymous):

/facepalm and I made another typo it seems: Yeah this is what I should have typed... 5*5*5*5*5*5*5*5*5*5*5*5+2*2*2*2*2*2*2*2*2*2

OpenStudy (anonymous):

I agree with @m_charron2

OpenStudy (anonymous):

yes any power of anything will not be prime but the problem is that there adding two powers... for example 2^2+3^2=13 see..

OpenStudy (anonymous):

Hmm, time to go back to the definitions then eh? http://en.wikipedia.org/wiki/Prime_number ^_^ (reading myself at the moment)

OpenStudy (anonymous):

but by the same token: 7^2+2^2 =16 so the question becomes do they make a prime or not

OpenStudy (anonymous):

7^2+2^2 = 53?

OpenStudy (anonymous):

oh derp.. math fail!, how about this the 7^1 +3^1 =10, there..

OpenStudy (anonymous):

Ah here we go, but I'll be honest and say that this is beyond my area of knowledge: http://en.wikipedia.org/wiki/AKS_primality_test Maybe somebody here could explain this? o_O

OpenStudy (anonymous):

Polynomial based primes

OpenStudy (anonymous):

yeah.. IDK

OpenStudy (turingtest):

@myininaya number theory help?

OpenStudy (anonymous):

@TuringTest Can you please explain the AKS_primality_test in layman's terms? I think that's probably the optimal answer to this question or any related to it that are sums of products (treating the variables of the polynomial as the product part)

OpenStudy (turingtest):

I wish I could, but that's why I tagged myininaya; I don't know number theory very well more possible helpers: @asnaseer @FoolForMath number theory help anyone?

OpenStudy (anonymous):

Guyssssss. It's a really basic solution. Just factorize it :P.